The diameter of the front wheel of a tricycle is 8 inches and of each of the two back wheels 3 inches. How many revolutions has each of the back wheels made when the front wheel has turned 1440 degrees?
A)2 2/3
B)4
c)10 2/3
D)24
E)96

Answer by stanbon(75887)

The diameter of the front wheel of a tricycle is 8 inches and of each of the two back wheels 3 inches. How many revolutions has each of the back wheels made when the front wheel has turned 1440 degrees?
----
# of front wheel revs = 1440/360 = 4
-----
distance traveled by front wheel: 4*(pi)*8 = 32(pi)inches
----
distance traveled by back wheel must also be 32(pi)inches
circumference of back wheel = 3(pi) inches
--
# of revs of back wheel = 32(pi)/3(pi) = 10 2/3
====================
Cheers,
Stan H.
====================






A)2 2/3
B)4
c)10 2/3
D)24
E)96